OpenStudy (anonymous):

Let f(x) = x-2/4 and g(x)= 2x^2+4. a. Find f(g(x)). b. Find g(f(x)).

OpenStudy (mathstudent55):

Do you know how to calculate f(1), for example, using f(x) you are given?

OpenStudy (anonymous):

No I have no clue where I start or what I am supposed to do. I am looking through my book right now to try to find an example.

OpenStudy (mathstudent55):

If you have a function, such as f(x) = 2x. Do you know how to do f(1)?

OpenStudy (anonymous):

No, I have no clue where I start or what I am supposed to do.

OpenStudy (anonymous):

f(g(x))= f(2x^2+4)= ??? g(f(X))= g(x-2/4)= ???

OpenStudy (mathstudent55):

Ok. Let's do it step by step. Let's say you have a function that is f(x) = 2x That means the same as y = 2x If you are asked to find f(1), all you need to do is substitute 1 in for x and calculate: f(x) = 2x Evaluate f(1). f(1) = 2(1) = 2, so f(1) = 2

OpenStudy (anonymous):

okay I understand that

OpenStudy (mathstudent55):

If you are asked to find f(t), then you replace x with t. Since in my example, f(x) = 2x, f(t) = 2t

OpenStudy (mathstudent55):

Now let's look at your problem. You have 2 given functions. One is called f and one is called g. Function f is defined as f(x) = (x - 2)/4, and fungtion g is defined as g(x) = 2x^2 + 4

OpenStudy (mathstudent55):

Look carefully at function f, f(x) = (x - 2)/4 If I were to ask you to find f(5), you would replace x with 5 and calculate. For you problem's f(x) = (x - 2)/4, f(5) is: f(5) = (5 - 2)/2 = 3/2

OpenStudy (anonymous):

f(g(x)) = f(2x^2+4) = x-2/4(2x^2+4) = ??? g(f(X)) = g(x-2/4) = 2x^2+4(x-2/4) = ???

OpenStudy (mathstudent55):

What your problem is asking is to find the value of f(x) evaluated at g(x). That is what f(g(x)) means. Instead of putting a number in for x, like putting 5 in for x to calculate f(5), they want you to put g(x) in for x.

OpenStudy (mathstudent55):

f(g(x)) = f(2x^2 + 4) = (2x^2 + 4 - 2)/4 = (2x^2 + 2)/4 = 2(x^2 + 1)/4 =(x^2 + 1)/2

OpenStudy (mathstudent55):